# MCPZoo: A Large-Scale Dataset of Runnable Model Context Protocol Servers for AI Agent

### v1

@@ Abstract
Model Context Protocol (MCP) enables agents to interact with external tools, yet empirical research on MCP is hindered by the lack of large-scale, accessible datasets. We present MCPZoo, the largest and most comprehensive dataset of MCP servers collected from multiple public sources, comprising 90,146 servers. MCPZoo includes over ten thousand server instances that have been deployed and verified as runnable and interactable, supporting realistic experimentation beyond static analysis. The dataset provides unified metadata and access interfaces, enabling systematic exploration and interaction without manual deployment effort. MCPZoo is released as an open and accessible resource to support research on MCP-based security analysis.
@@ 1 Introduction
By providing a standardized, “unrestricted, and scalable” dataset of runnable agents, we aim to enable the research community to move beyond small-scale testing and conduct comprehensive, reproducible analyses on the security and capability of the AI agent ecosystem.
@@
  • The largest and most comprehensive dataset of MCP servers currently available, with 90,146 MCP servers.
@@ 2.1 Construct MCP Zoo
Sources. We collected MCP server projects from six public sources: MCP World [mcpworld], MCP Store [mcpstore], MCP Servers Repository [mcprepository], AIbase MCP [aibasemcp], Mcp.so [mcpso], and Pulse MCP [pulsemcp]. Each source provides an independently curated directory of MCP servers, reflecting different community and platform perspectives.
@@ 2.2 Make the Zoo Alive
Building Docker Image. To make servers runnable at scale, we employ an automated build agent to convert the MCP server source code into executable Docker images. The agent performs context-aware analysis of code and dependencies to infer runtime settings and generate Docker configurations. When build failures occur, it iteratively analyzes error signals and regenerates the Docker files with targeted adjustments until the image builds successfully or a retry limit is reached.
Liveness Check. A successfully built container does not necessarily indicate a runnable MCP server. We therefore perform a standardized liveness check to validate whether a server is practically usable in the MCP setting. This check consists of two complementary aspects. The first aspect examines execution capability, accounting for common failure cases such as missing configurations or hardcoded paths. The second aspect examines protocol-level functionality through basic interactions, including retrieving declared tool lists. Only servers that pass both execution and interaction checks are considered alive in MCPZoo.
@@ 3.1 Overview
MCPZoo is a large-scale, continuously growing dataset of Model Context Protocol (MCP) servers collected from multiple public sources. At the time of writing, MCPZoo contains 90,146 distinct MCP servers, whose source code collectively amounts to 377 GB, among which 14,206 servers are verified to be runnable. Table I summarizes the composition of MCPZoo by source, reporting the number of collected servers from each of the platforms.
TABLE I: Statistics of MCP Servers from Six Public Sources
@@
MMCP World 31,048 34.44%
MCP Store 25,695 28.50%
MCP Repository 14,341 15.91%
AIbase MCP 11,120 12.33%
Mcp.so 6,772 7.51%
Pulse MCP 1,170 1.30%
Total 90,146 100.00%

### v2

@@ Abstract
Model Context Protocol (MCP) enables agents to interact with external tools, yet empirical research on MCP is hindered by the lack of large-scale, accessible datasets. We present MCPZoo, the largest and most comprehensive dataset of MCP servers collected from multiple public sources, comprising 95,142 servers. MCPZoo includes over ten thousand server instances that have been deployed and verified as runnable and interactable, supporting realistic experimentation beyond static analysis. The dataset provides unified metadata and access interfaces, enabling systematic exploration and interaction without manual deployment effort. MCPZoo is released as an open and accessible resource to support research on MCP-based security analysis.
@@ 1 Introduction
By providing a standardized, unrestricted, and scalable dataset of runnable agents, we aim to enable the research community to move beyond small-scale testing and conduct comprehensive, reproducible analyses on the security and capability of the AI agent ecosystem.
@@
  • The largest and most comprehensive dataset of MCP servers currently available, with 95,142 MCP servers.
@@ 2.1 Construct MCP Zoo
Sources. We collected MCP server projects from seven public sources: MCP World [mcpworld], MCP Store [mcpstore], MCP Servers Repository [mcprepository], AIbase MCP [aibasemcp], Mcp.so [mcpso], MCP Market [mcpmarket], and Pulse MCP [pulsemcp]. Each source provides an independently curated directory of MCP servers, reflecting different community and platform perspectives.
@@ 2.2 Make the Zoo Alive
Building Docker Image. To make servers runnable at scale, we employ an automated build agent to convert the MCP server source code into executable Docker images. The agent performs context-aware analysis of code and dependencies to infer runtime settings and generate Docker configurations. When build failures occur, it iteratively analyzes error logs and regenerates the Docker files with targeted adjustments until the image builds successfully or a retry limit is reached.
Liveness Check. A successfully built container does not necessarily indicate a runnable MCP server. We therefore perform a standardized liveness check to validate whether a server is practically usable in the MCP setting. We perform a two-step check to guarantee the liveness. First, we examine the containers’ execution capability, accounting for common failure cases such as missing configurations or hard-coded paths. Then we examine the server containers’ protocol-level functionality through basic interactions, including retrieving declared tool lists. Only servers that pass both execution and interaction checks are considered alive in MCPZoo.
@@ 3.1 Overview
MCPZoo is a large-scale, continuously growing dataset of Model Context Protocol (MCP) servers collected from multiple public sources. Up to December 17th, 2025, MCPZoo contains 95,142 distinct MCP servers, whose source code collectively amounts to 377 GB, among which 14,206 servers are verified to be runnable. Table I summarizes the composition of MCPZoo by source, reporting the number of collected servers from each of the platforms.
TABLE I: Statistics of MCP Servers from Seven Public Sources
@@
MCP World 31,048 32.63%
MCP Store 25,695 27.00%
MCP Repository 14,341 15.07%
AIbase MCP 11,120 11.69%
Mcp.so 6,772 7.12%
MCP Market 4,996 5.25%
Pulse MCP 1,170 1.23%
Total 95,142 100.0%